OpenStudy (anonymous):

Applied Calculus 1 question. How would you go about finding the limit of the following function without using a calculator?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5} \frac{ \sqrt{x+4}-3 }{ x-5 }\]

zepdrix (zepdrix):

Hey Mojo :) You would start by applying an algebraic trick, multiplying the top and bottom by the `conjugate` of the numerator.

zepdrix (zepdrix):

\[\large\rm \lim_{x\to5}\frac{\sqrt{x+4}-3}{x-5}\color{royalblue}{\left(\frac{\sqrt{x+4}+3}{\sqrt{x+4}+3}\right)}\]

OpenStudy (anonymous):

yup did that got \[\lim_{x \rightarrow 5}\frac{ x+1 }{ (x-5)(\sqrt{x+4}-3) }\]

OpenStudy (anonymous):

sorry mean to have +3 not negative

zepdrix (zepdrix):

Hmmm that numerator looks a big messed up :O

OpenStudy (anonymous):

oooh i see what id id wrong lol

zepdrix (zepdrix):

Recall that when you multiply conjugates, you get the difference of `squares`. You have to square that 3! :)

OpenStudy (anonymous):

thanks for pointing that out

OpenStudy (anonymous):

yay the got an answer of 1/9

OpenStudy (anonymous):

thanks again